### 10 Mar 2026 v1

# The Hofstadter consecutive-sum sequence omits infinitely many positive integers

Abstract.
Let $(a_{n})_{n\geq 1}$ be the greedy self-generating sequence defined by $a_{1}=1$, $a_{2}=2$, and, for $k\geq 3$, by taking $a_{k}$ to be the least integer greater than $a_{k-1}$ that can be written as a sum of at least two consecutive earlier terms. Hofstadter asked about the asymptotic behavior of this sequence. In this paper we prove that
$n+\omega(1)\leq a_{n}\ll n^{4175/2506+o(1)}.$
In particular, $(a_{n})_{n\geq 1}$ omits infinitely many positive integers, thereby settling a conjecture from the OEIS entry A005243.
Key words and phrases:
self-generating sequence, consecutive-term sums
2020 Mathematics Subject Classification:
Primary 11B83; Secondary 11D61
1. Introduction
A self-generating greedy process, often discussed under the name Hofstadter sequences (see, for instance, [9] or [6, E31]), starts from a short initial list and then repeatedly adjoins sums of consecutive earlier terms. In its classical form, the seed is $(1,2)$, and one generates a strictly increasing sequence $(a_{n})_{n\geq 1}$ by setting $a_{1}=1$, $a_{2}=2$, and, for $k\geq 3$, taking $a_{k}$ to be the least integer greater than $a_{k-1}$ that can be written as a sum of at least two consecutive earlier terms:
$a_{k}=\sum_{i=p}^{q}a_{i}\qquad\text{for some }1\leq p\leq q\leq k-1,\ \ q-p\geq 1.$ (1.1)
This is A005243 in the OEIS [8]. The asymptotic behavior of $(a_{n})_{n\geq 1}$ was asked about by Hofstadter and has been recorded in several places, including [4, p. 71], [5, p. 83], [6, E31], and as Problem #423 on Bloom’s Erdős Problems website [2].
Problem 1.1 ([4, 5, 6]).
What is the asymptotic behavior of the sequence (1.1)?
Numerically, the sequence appears to contain “most” integers. For orientation, the sequence (1.1) begins
$1,2,3,5,6,8,10,11,14,16,17,18,19,21,22,24,25,29,30,32,\dots,$
and hence it already omits $4,7,9,12,13,15,20,\dots$. At the same time, the COMMENTS section of A005243 records the following conjecture.
Conjecture 1.2 ([8]).
The sequence (1.1) omits infinitely many positive integers.
In this paper we prove Conjecture 1.2. In fact, our argument applies to a much broader class of related greedy sequences obtained from an arbitrary finite initial seed.
Definition 1.3.
Let $s\geq 2$ and let $\mathbf{u}=(u_{1},\dots,u_{s})$ be a seed of length $s$ consisting of positive integers (not necessarily distinct or ordered). Define a sequence of positive integers $(a_{n}^{\mathbf{u}})_{n\geq 1}$ by
$a_{n}^{\mathbf{u}}:=u_{n}\qquad(1\leq n\leq s),$
and for $n\geq s+1$ let $a_{n}^{\mathbf{u}}$ be the least integer greater than $a_{n-1}^{\mathbf{u}}$ that can be written as a sum of at least two consecutive earlier terms:
$a_{n}^{\mathbf{u}}=\sum_{k=p}^{q}a_{k}^{\mathbf{u}}\qquad\text{for some }1\leq p\leq q\leq n-1,\ \ q-p\geq 1.$
We also set
$b_{n}^{\mathbf{u}}:=a_{n}^{\mathbf{u}}-n\qquad(n\geq 1).$
This definition is well defined: for each $n\geq s+1$, the sum $a_{n-2}^{\mathbf{u}}+a_{n-1}^{\mathbf{u}}$ is a valid candidate for $a_{n}^{\mathbf{u}}$, and it is strictly greater than $a_{n-1}^{\mathbf{u}}$ because all terms are positive. Therefore the minimum in the definition of $a_{n}^{\mathbf{u}}$ always exists. In particular, we have $a_{n+1}^{\mathbf{u}}>a_{n}^{\mathbf{u}}$ for all $n\geq s$.
When $\mathbf{u}=(1,2)$, the sequence $(a_{n}^{\mathbf{u}})_{n\geq 1}$ coincides with the classical sequence (1.1).
Theorem 1.4.
For every integer $s\geq 2$ and every seed $\mathbf{u}$ of length $s$ consisting of positive integers, the sequence $(b_{n}^{\mathbf{u}})_{n\geq 1}$ is eventually nondecreasing; indeed, it is nondecreasing for all $n\geq s$. Moreover, $(b_{n}^{\mathbf{u}})_{n\geq 1}$ is unbounded.
The quantity $b_{n}^{\mathbf{u}}=a_{n}^{\mathbf{u}}-n$ measures the excess of the current term over its index. For all sufficiently large $n$, it gives a lower bound for the number of omitted positive integers in $1,a_{n}^{\mathbf{u}}]$. As a consequence, we obtain the following corollary.
Corollary 1.5.
For every integer $s\geq 2$ and every seed $\mathbf{u}$ of length $s$ consisting of positive integers,
$a_{n}^{\mathbf{u}}=n+\omega(1).$
In particular, the set $\{a_{n}^{\mathbf{u}}:n\geq 1\}$ omits infinitely many positive integers.
In particular, taking $\mathbf{u}=(1,2)$ in Corollary 1.5 proves Conjecture 1.2.
For the classical sequence (1.1), we also obtain a quantitative upper bound. In fact, the same argument extends to every fixed positive seed; see Remark 4.1.
Theorem 1.6.
For every $\varepsilon>0$ there exists $C_{\varepsilon}>0$ such that, for all $n\geq 1$,
$a_{n}\leq C_{\varepsilon}\,n^{4175/2506+\varepsilon}.$
The proof of Theorem 1.6 combines the greedy structure of the sequence with a recent lower bound for difference sets of finite convex sets. Thus, besides proving that the sequence omits infinitely many positive integers, we also obtain a first polynomial upper bound toward Problem 1.1.
Notation
We use Vinogradov’s asymptotic notation. For functions $f=f(n)$ and $g=g(n)$, we write $f=O(g)$, $f\ll g$, or $g\gg f$ to mean that there exists a constant $C>0$ such that $|f(n)|\leq Cg(n)$ for all sufficiently large $n$. If the implicit constant is allowed to depend on one or more parameters $z_{1},\dots,z_{r}$, we indicate this by writing $f\ll_{z_{1},\dots,z_{r}}g$, $g\gg_{z_{1},\dots,z_{r}}f$ or $f=O_{z_{1},\dots,z_{r}}(g)$. We write $f\asymp g$ or $f=\Theta(g)$ to mean that $f\ll g$ and $g\ll f$, and write $f=o(g)$ to mean $f(n)/g(n)\to 0$ as $n\to\infty$. We write $f=\omega(g)$ to mean that $g=o(f)$, equivalently, $|f(n)|/g(n)\to\infty$ when $n\to\infty$, assuming $g(n)>0$ for all sufficiently large $n$. In particular, $\omega(1)$ denotes any quantity tending to infinity as $n\to\infty$.
Paper organization
Section 2 collects the preliminary ingredients, including the characterization of integers that can be written as sums of at least two consecutive positive integers and a Diophantine finiteness consequence of a result of Schinzel and Tijdeman. In Section 3 we prove Theorem 1.4 and Corollary 1.5. Section 4 is devoted to the proof of the polynomial upper bound in Theorem 1.6. Finally, Section 5 contains concluding remarks and numerical observations.
2. Preliminaries
We start with a standard characterization of sums of consecutive integers.
Lemma 2.1.
A positive integer $N$ can be written as a sum of at least two consecutive positive integers if and only if $N$ is not a power of $2$.
Proof.
Suppose $N=x+(x+1)+\cdots+(x+\ell-1)$ with $\ell\geq 2$ and $x\geq 1$. Then
$2N=\ell(2x+\ell-1).$
If $\ell$ is odd, then $\ell\mid 2N$ implies $\ell\mid N$, so $N$ has an odd divisor >1>1. If $\ell$ is even, then $2x+\ell-1$ is odd and divides $2N$, hence divides $N$; again $N$ has an odd divisor >1>1. In either case $N$ cannot be a power of $2$.
Conversely, if $N$ is not a power of $2$, then $N$ has an odd divisor $d>1$, say $N=dm$. If $m\geq(d+1)/2$, then
$N=\Bigl(m-\frac{d-1}{2}\Bigr)+\Bigl(m-\frac{d-1}{2}+1\Bigr)+\cdots+\Bigl(m+\frac{d-1}{2}\Bigr)$
is a sum of $d$ consecutive positive integers. If $m<(d+1)/2$, then
$N=\Bigl(\frac{d+1}{2}-m\Bigr)+\Bigl(\frac{d+1}{2}-m+1\Bigr)+\cdots+\Bigl(\frac{d+1}{2}+m-1\Bigr)$
is a sum of $2m\geq 2$ consecutive positive integers. ∎
Next we record the Diophantine ingredient. We shall use the following consequence of a result of Schinzel and Tijdeman [7, Corollary 1].
Lemma 2.2 ([7]).
Let $P(x)\in\mathbb{Q}[x]$ be a polynomial with at least two simple zeros. Then the equation
$y^{k}=P(x),\qquad x,y\in\mathbb{Z},\ \ |y|>1,$
has only finitely many integer solutions $(x,y,k)$ with $k>2$.
Corollary 2.3.
Fix $E\in\mathbb{Z}$. The equation
$v^{2}+v+E=2^{k}$
has only finitely many integer solutions $(v,k)\in\mathbb{Z}\times\mathbb{Z}$ with $k>2$.
Proof.
Apply Lemma 2.2 with $y=2$ and $P(v)=v^{2}+v+E$. Since $E\in\mathbb{Z}$, the discriminant of $P$ is $1-4E\neq 0$; hence $P$ has two distinct (and therefore simple) zeros, so the lemma applies. ∎
3. A qualitative lower bound and infinitely many omissions
Throughout this section, we fix a seed $\mathbf{u}$ of length $s\geq 2$, and write
$a_{n}:=a_{n}^{\mathbf{u}},\qquad b_{n}:=b_{n}^{\mathbf{u}}\qquad(n\geq 1).$
Our goal is to show that $(b_{n})$ is unbounded. The first step is a simple monotonicity observation, which will later force eventual linearity if $(b_{n})$ were bounded.
3.1. Monotonicity and eventual linearity
Lemma 3.1.
The sequence $(b_{n})_{n\geq s}$ is nondecreasing.
Proof.
For every $n\geq s$, the greedy construction gives $a_{n+1}>a_{n}$, hence $a_{n+1}\geq a_{n}+1$. Therefore, for $n\geq s$,
$b_{n+1}-b_{n}=(a_{n+1}-(n+1))-(a_{n}-n)=a_{n+1}-a_{n}-1\geq 0.\qed$
Lemma 3.2.
If $(b_{n})_{n\geq s}$ is bounded above, then there exist integers $n_{0}\geq s$ and $B$ such that
$a_{n}=n+B\qquad\text{for all }n\geq n_{0}.$
Proof.
By Lemma 3.1, the sequence $(b_{n})_{n\geq s}$ is a nondecreasing sequence of integers. If $(b_{n})_{n\geq s}$ is bounded above, then $(b_{n})_{n\geq s}$ must be eventually constant: there exist $n_{0}\geq s$ and an integer $B$ such that $b_{n}=B$ for all $n\geq n_{0}$. Hence $a_{n}=n+b_{n}=n+B$ for all $n\geq n_{0}$. ∎
From now on we assume, for contradiction, that $(b_{n})_{n\geq s}$ is bounded above and hence that
$a_{n}=n+B\qquad(n\geq n_{0})$ (3.1)
for some integers $n_{0}\geq s$ and $B$.
3.2. Consecutive-block decompositions beyond the linear tail
Set $T:=n_{0}+B$ and
$S_{n_{0}-1}:=\sum_{i=1}^{n_{0}-1}a_{i},\qquad\mathcal{C}:=\Bigl\{\sum_{k=p}^{n_{0}-1}a_{k}:\ 1\leq p\leq n_{0}\Bigr\}.$
(Here $p=n_{0}$ contributes the empty sum $0\in\mathcal{C}$.) The next lemma isolates the only two ways a large term can be written as a sum of consecutive earlier terms once the tail is linear.
Lemma 3.3.
Assume (3.1). Let $t>S_{n_{0}-1}$, and suppose that
$t=\sum_{k=p}^{q}a_{k}\qquad\text{for some }1\leq p<q.$
Then $q\geq n_{0}$, and exactly one of the following holds:
  • (i)
    $p\geq n_{0}$, in which case $t=\sum_{j=p+B}^{q+B}j$;
  • (ii)
    $p<n_{0}\leq q$, in which case $t=C+\sum_{j=T}^{q+B}j$ for some $C\in\mathcal{C}$.
Proof.
If $q\leq n_{0}-1$, then
$t=\sum_{k=p}^{q}a_{k}\leq\sum_{k=1}^{n_{0}-1}a_{k}=S_{n_{0}-1},$
contrary to the assumption $t>S_{n_{0}-1}$. Hence $q\geq n_{0}$.
If $p\geq n_{0}$, then by (3.1),
$t=\sum_{k=p}^{q}(k+B)=\sum_{j=p+B}^{q+B}j,$
which is (i).
If $p<n_{0}\leq q$, then splitting the sum at $n_{0}$ gives
$t=\sum_{k=p}^{n_{0}-1}a_{k}+\sum_{k=n_{0}}^{q}a_{k}=C+\sum_{k=n_{0}}^{q}(k+B)=C+\sum_{j=T}^{q+B}j$
for some $C\in\mathcal{C}$, which is (ii).
The two alternatives are mutually exclusive, since they impose opposite inequalities on $p$. ∎
3.3. Powers of two force a quadratic exponential equation
The linear tail (3.1) implies that sufficiently large powers of two appear among the $a_{n}$. We now analyze the corresponding representations of these powers of two as sums of consecutive earlier terms.
Lemma 3.4.
Assume (3.1). For every integer $r$ with
$2^{r}\geq\max\{T,\ S_{n_{0}-1}+1,\ B+s+1\},$ (3.2)
there exist $C\in\mathcal{C}$ and an integer $v\geq T$ such that
$2^{r+1}=v^{2}+v+E_{C},\qquad E_{C}:=2C-(T-1)T.$ (3.3)
Proof.
Fix $r$ satisfying the inequality (3.2) and set $n=2^{r}-B$. Then $n\geq n_{0}$ and (3.1) gives $a_{n}=n+B=2^{r}$. Moreover, $n\geq s+1$, so by construction of the sequence there exist indices $1\leq p\leq q\leq n-1$ with $q-p\geq 1$ and
$2^{r}=a_{n}=\sum_{k=p}^{q}a_{k}.$
Since $2^{r}>S_{n_{0}-1}$, Lemma 3.3 applies. The alternative (i) is impossible: if $2^{r}=\sum_{j=p+B}^{q+B}j$ with $q\geq p+1$, then $p\geq n_{0}$ implies
$p+B\geq n_{0}+B=T=a_{n_{0}}>0.$
Hence $2^{r}$ would be a sum of at least two consecutive positive integers, contradicting Lemma 2.1. Therefore we are in case (ii). Thus for some $C\in\mathcal{C}$ and some $v:=q+B\geq T$ we have
$2^{r}=C+\sum_{j=T}^{v}j=C+\frac{v(v+1)-(T-1)T}{2}.$
Multiplying by $2$ yields (3.3). ∎
3.4. Completion of the proof
Proof of Theorem 1.4.
Lemma 3.1 shows that $(b_{n})$ is nondecreasing for all $n\geq s$. It remains to prove that it is unbounded.
Assume, for contradiction, that $(b_{n})$ is bounded above. Then (3.1) holds by Lemma 3.2. Lemma 3.4 shows that for every sufficiently large $r$ we can find $C\in\mathcal{C}$ and $v\geq T$ such that (3.3) holds. The set
$\mathcal{E}:=\{E_{C}:\ C\in\mathcal{C}\}$
is finite. Hence by the pigeonhole principle there exists some $E\in\mathcal{E}$ for which the equation
$v^{2}+v+E=2^{k}$
has infinitely many integer solutions $(v,k)\in\mathbb{Z}\times\mathbb{Z}$ with $k>2$. This contradicts Corollary 2.3. Therefore $(b_{n})$ is unbounded, as claimed. ∎
Proof of Corollary 1.5.
By Theorem 1.4, the sequence $(b_{n})$ is nondecreasing for all $n\geq s$ and unbounded, hence $b_{n}\to\infty$. Therefore
$a_{n}=n+b_{n}=n+\omega(1).$
Let $M:=\max_{1\leq i\leq s}a_{i}$. For all sufficiently large $n$, we have $n\geq s$ and $a_{n}\geq M$. Hence every value attained by the sequence in $1,a_{n}]$ occurs among the first $n$ terms: indeed, the initial seed values are all $\leq M\leq a_{n}$, while $a_{m}>a_{n}$ for every $m>n$ because the sequence is strictly increasing from index $s$ onward. The set of distinct values among $a_{1},\dots,a_{n}$ therefore has cardinality at most $n$, so the number of positive integers in $1,a_{n}]$ omitted by the sequence is at least
$a_{n}-n=b_{n}.$
Since $b_{n}\to\infty$, the number of omitted integers up to $a_{n}$ is unbounded. In particular, the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers. ∎
4. A quantitative upper bound for the classical sequence
In this section we restrict attention to the classical sequence (1.1) and prove the bound
$a_{n}\ll_{\varepsilon}n^{4175/2506+\varepsilon}.$
Remark 4.1.
For every fixed seed $\mathbf{u}$ of length $s\geq 2$ consisting of positive integers, the same exponent can also be obtained for $a_{n}^{\mathbf{u}}$, namely $a_{n}^{\mathbf{u}}\ll_{\mathbf{u},\varepsilon}n^{4175/2506+\varepsilon}$. Indeed, by Definition 1.3, the sequence $(a_{n}^{\mathbf{u}})_{n\geq s}$ is strictly increasing, so one may apply the argument of the present section to the tail
$a_{s}^{\mathbf{u}},a_{s+1}^{\mathbf{u}},a_{s+2}^{\mathbf{u}},\dots$
after making the obvious notational changes in the definitions of the relevant prefix sums and representable integers. Since this introduces additional bookkeeping but no essentially new idea, we present the details only for the classical seed $(1,2)$.
4.1. Greedy enumeration of the representable integers
We begin by isolating the set of integers that can be represented as sums of at least two consecutive terms. The key point is that the greedy construction does not merely produce some such integers: it lists all of them in increasing order.
Define the partial sums
$s_{0}:=0,\qquad s_{m}:=\sum_{i=1}^{m}a_{i}\quad(m\geq 1),$
and the prefix-sum set
$\widetilde{S}_{m}:=\{s_{0},s_{1},\dots,s_{m}\}.$
For $m\geq 2$ let
$\mathcal{R}_{m}:=\Bigl\{\sum_{i=p}^{q}a_{i}:\ 1\leq p<q\leq m\Bigr\}\subset\mathbb{N},\qquad\mathcal{R}:=\bigcup_{m\geq 2}\mathcal{R}_{m}.$
Thus $\mathcal{R}$ is the set of all integers representable as a sum of at least two consecutive terms of the sequence.
Lemma 4.2.
We have $\{a_{k}:k\geq 3\}=\mathcal{R}$, and $(a_{3},a_{4},\dots)$ is the increasing enumeration of $\mathcal{R}$. In particular, for every $X\geq 1$,
$(\mathcal{R}\cap[1,X])\ \geq\ n\quad\Longrightarrow\quad a_{n+2}\leq X.$ (4.1)
Proof.
First note that $a_{k}\in\mathcal{R}_{k-1}\subset\mathcal{R}$ for every $k\geq 3$ by definition, so $\{a_{k}:k\geq 3\}\subseteq\mathcal{R}$.
Fix $k\geq 3$. We claim that there is no element of $\mathcal{R}$ strictly between $a_{k-1}$ and $a_{k}$. Indeed, suppose for contradiction that there exists $x\in\mathcal{R}$ with
$a_{k-1}<x<a_{k}.$
Choose a representation $x=\sum_{i=p}^{q}a_{i}$ with $p<q$. Since all terms are positive, every summand satisfies $a_{i}\leq x<a_{k}$. As $(a_{i})$ is strictly increasing, this forces $q\leq k-1$, so in fact $x\in\mathcal{R}_{k-1}$. But then $x$ is a representable integer using the first k1k-1 terms and $x>a_{k-1}$, contradicting the greedy choice of $a_{k}$ as the least such representable integer. This proves the claim.
It remains to show the reverse inclusion $\mathcal{R}\subseteq\{a_{k}:k\geq 3\}$. Let $x\in\mathcal{R}$. Since $(a_{k})$ is strictly increasing with $a_{k}\geq a_{k-1}+1$, we have $a_{k}\to\infty$, so there exists a minimal $k\geq 3$ such that $a_{k}\geq x$. Then $a_{k-1}<x\leq a_{k}$. If $x<a_{k}$, this would place $x$ in the interval $(a_{k-1},a_{k})$, contradicting the claim proved above that $\mathcal{R}$ contains no element strictly between $a_{k-1}$ and $a_{k}$. Hence $x=a_{k}$, proving $\mathcal{R}\subseteq\{a_{k}:k\geq 3\}$. Consequently, for each $k\geq 3$, the number $a_{k}$ is the smallest element of $\mathcal{R}$ that is greater than $a_{k-1}$. Therefore $(a_{3},a_{4},\dots)$ lists the elements of $\mathcal{R}$ in strictly increasing order, i.e. it is the increasing enumeration of $\mathcal{R}$.
Finally, (4.1) follows immediately: if $(\mathcal{R}\cap[1,X])\geq n$ then the $n$-th element of $\mathcal{R}$ is $\leq X$, hence $a_{n+2}\leq X$. ∎
The preceding lemma reduces the problem of bounding $a_{n}$ to that of producing many representable integers below a controlled threshold. The next observation makes this reduction quantitative at a finite stage.
Lemma 4.3.
For every $m\geq 2$, we have $\mathcal{R}_{m}\subseteq\mathcal{R}\cap[1,s_{m}]$. In particular, if $|\mathcal{R}_{m}|\geq n$ then
$a_{n+2}\leq s_{m}.$ (4.2)
Proof.
Every element of $\mathcal{R}_{m}$ is a sum of at least two consecutive terms among $a_{1},\dots,a_{m}$, hence belongs to $\mathcal{R}$. Such a sum is clearly at most $a_{1}+\cdots+a_{m}=s_{m}$, so $\mathcal{R}_{m}\subseteq\mathcal{R}\cap[1,s_{m}]$.
If $|\mathcal{R}_{m}|\geq n$, then $(\mathcal{R}\cap[1,s_{m}])\geq n$, and (4.2) follows from Lemma 4.2 and (4.1). ∎
We now turn to the task of estimating $|\mathcal{R}_{m}|$. The key idea is to rewrite sums of consecutive terms as differences of prefix sums, thereby bringing in additive-combinatorial results on convex sets.
4.2. From consecutive sums to a convex difference set
We recall the standard additive-combinatorial notion of convexity. A finite set $Y=\{y_{1}<y_{2}<\cdots<y_{n}\}\subset\mathbb{R}$ is called convex if its consecutive differences form a strictly increasing sequence, that is,
$y_{i+1}-y_{i}<y_{i+2}-y_{i+1}\qquad(1\leq i\leq n-2).$
We now observe that the prefix-sum set $\widetilde{S}_{m}$ is convex in this sense.
Lemma 4.4.
For each $m\geq 2$, the set $\widetilde{S}_{m}=\{s_{0}<s_{1}<\cdots<s_{m}\}$ is convex, i.e.
$s_{i}-s_{i-1}<s_{i+1}-s_{i}\qquad(1\leq i\leq m-1).$
Proof.
We have $s_{i}-s_{i-1}=a_{i}$ for all $i\geq 1$, and $(a_{i})$ is strictly increasing by construction. ∎
Since every sum of consecutive terms is a difference of two prefix sums, the difference set $\widetilde{S}_{m}-\widetilde{S}_{m}$ provides a natural ambient set in which $\mathcal{R}_{m}$ sits. The next lemma makes this relation explicit.
Lemma 4.5.
Let $D_{m}:=\widetilde{S}_{m}-\widetilde{S}_{m}=\{x-y:\ x,y\in\widetilde{S}_{m}\}$. Then
$|\mathcal{R}_{m}|\geq\frac{|D_{m}|-1}{2}-m.$
Proof.
Let
$D_{m}^{+}:=\{d\in D_{m}:\ d>0\}.$
Since $D_{m}$ is symmetric and contains 0, we have
$|D_{m}^{+}|=\frac{|D_{m}|-1}{2}.$
Every $d\in D_{m}^{+}$ can be written as
$d=s_{j}-s_{i}=\sum_{t=i+1}^{j}a_{t}\qquad(0\leq i<j\leq m).$
If ji=1j-i=1, then $d=a_{j}\in\{a_{1},\dots,a_{m}\}$. If $j-i\geq 2$, then $d\in\mathcal{R}_{m}$. Therefore
$D_{m}^{+}\subseteq\{a_{1},\dots,a_{m}\}\cup\mathcal{R}_{m}.$
Hence
$\frac{|D_{m}|-1}{2}=|D_{m}^{+}|\leq m+|\mathcal{R}_{m}|,$
which rearranges to
$|\mathcal{R}_{m}|\geq\frac{|D_{m}|-1}{2}-m.\qed$
4.3. Bloom’s convex difference-set bound and the basic recursion
We now combine the convexity of the prefix-sum set with a recent difference-set estimate of Bloom. This will yield a lower bound for $|\mathcal{R}_{m}|$, and hence, via the greedy enumeration established above, a recursive upper bound for $a_{n}$.
The key additive-combinatorial input is the following difference-set bound for finite convex sets, due to Bloom [1, Theorem 2].
Theorem 4.6 ([1]).
For every $\eta>0$ there exists $c_{\eta}>0$ such that for every finite convex set $A\subset\mathbb{R}$,
$|A-A|\geq c_{\eta}\,|A|^{6681/4175-\eta}.$
Applying this estimate to the convex set $\widetilde{S}_{m}$, and then using the relation between $\widetilde{S}_{m}-\widetilde{S}_{m}$ and $\mathcal{R}_{m}$ established in Lemma 4.5, we obtain the following lower bound for the number of representable integers up to stage $m$.
Proposition 4.7.
Fix $0<\eta<2506/4175$ and set $\delta=6681/4175-\eta$. Then there exist constants $m_{0}\in\mathbb{N}$ and $c>0$ (depending only on $\eta$) such that for all $m\geq m_{0}$,
$|\mathcal{R}_{m}|\geq c\,m^{\delta}.$
Proof.
By Lemma 4.4, the set $\widetilde{S}_{m}$ is convex and $|\widetilde{S}_{m}|=m+1$. Applying Theorem 4.6 to $A=\widetilde{S}_{m}$ gives
$|D_{m}|=|\widetilde{S}_{m}-\widetilde{S}_{m}|\geq c_{\eta}\,(m+1)^{\delta}.$
Insert this into Lemma 4.5:
$|\mathcal{R}_{m}|\geq\frac{c_{\eta}(m+1)^{\delta}-1}{2}-m.$
For all sufficiently large $m$ the right-hand side is $\geq(c_{\eta}/4)\,m^{\delta}$, which proves this proposition with suitable choices of $m_{0}$ and $c$. ∎
The proposition shows that $\mathcal{R}_{m}$ grows polynomially with exponent $\delta>1$. We now feed this counting information back into Lemma 4.3 to obtain the basic recursion for $(a_{n})$.
Lemma 4.8.
Fix $0<\eta<2506/4175$ and let $\delta=6681/4175-\eta>1$. Then there exist constants $C_{1}\geq 2$ and $N_{0}$ (depending only on $\eta$) such that for all $n\geq N_{0}$,
$a_{n}\ \leq\ C_{1}\,n^{1/\delta}\,a_{\lceil C_{1}\,n^{1/\delta}\rceil}.$ (4.3)
Proof.
Let $c,m_{0}$ be as in Proposition 4.7. For $N$ large, set
$m_{1}:=\left\lceil(N/c)^{1/\delta}\right\rceil.$
Then $m_{1}\geq m_{0}$ and $cm_{1}^{\delta}\geq N$, so $|\mathcal{R}_{m_{1}}|\geq N$. By Lemma 4.3,
$a_{N+2}\leq s_{m_{1}}.$
Since $a_{i}\leq a_{m_{1}}$ for $1\leq i\leq m_{1}$, we have $s_{m_{1}}\leq m_{1}a_{m_{1}}$, hence
$a_{N+2}\leq m_{1}a_{m_{1}}.$
Moreover, for $N$ large enough,
$m_{1}=\left\lceil(N/c)^{1/\delta}\right\rceil\leq 2c^{-1/\delta}N^{1/\delta}.$
Set
$C_{1}:=\max\{2,\,2c^{-1/\delta}\}.$
Using the monotonicity of $(a_{n})$, we obtain
$a_{N+2}\leq C_{1}N^{1/\delta}a_{\lceil C_{1}N^{1/\delta}\rceil}$
for all sufficiently large $N$.
Now replace $N$ by n2n-2. For all sufficiently large $n$,
$a_{n}\leq C_{1}(n-2)^{1/\delta}a_{\lceil C_{1}(n-2)^{1/\delta}\rceil}\leq C_{1}n^{1/\delta}a_{\lceil C_{1}n^{1/\delta}\rceil}.$
This is (4.3) after renaming the lower threshold as $N_{0}$. ∎
4.4. Iteration and the final exponent
The basic recursion from Lemma 4.8 can now be iterated. The next lemma shows that any nondecreasing sequence satisfying such a recursion must obey a polynomial upper bound, up to a logarithmic factor.
Lemma 4.9.
Let $(a_{n})$ be nondecreasing and suppose there exist constants $C_{1}\geq 2$, $\alpha\in(0,1)$ and $N_{0}$ such that
$a_{n}\leq C_{1}\,n^{\alpha}\,a_{\lceil C_{1}n^{\alpha}\rceil}\qquad(n\geq N_{0}).$
Then there exist constants $B,C_{2}>0$ such that for all $n\geq 3$,
$a_{n}\leq C_{2}\,n^{\alpha/(1-\alpha)}\,(\log n)^{B}.$
Proof.
Enlarge $N_{0}$ if necessary so that
$\lceil C_{1}x^{\alpha}\rceil\leq 2C_{1}x^{\alpha}\leq x-1\qquad(x\geq N_{0}),$
and also
$\frac{\log(2C_{1})}{1-\alpha}\leq\frac{1}{2}\log N_{0}.$
Fix $n>N_{0}$, and define
$\widetilde{n}_{0}:=n,\qquad\widetilde{n}_{j+1}:=\lceil C_{1}\widetilde{n}_{j}^{\alpha}\rceil\quad(j\geq 0)$
for as long as $\widetilde{n}_{j}>N_{0}$. Since $\widetilde{n}_{j+1}\leq\widetilde{n}_{j}-1$ whenever $\widetilde{n}_{j}\geq N_{0}$, the positive integer sequence $(\widetilde{n}_{j})$ strictly decreases until it first reaches the interval $1,N_{0}]$. Hence there exists a minimal $t=t(n)$ such that $\widetilde{n}_{t}\leq N_{0}$.
Repeatedly applying the assumed recursion gives
$a_{n}\ \leq\ \Bigl(\prod_{j=0}^{t-1}C_{1}\widetilde{n}_{j}^{\alpha}\Bigr)\,a_{\widetilde{n}_{t}}.$
Since $\widetilde{n}_{t}\leq N_{0}$, we have
$a_{\widetilde{n}_{t}}\leq C_{0}:=\max_{1\leq r\leq N_{0}}a_{r}.$
Also, from $\widetilde{n}_{j+1}\leq 2C_{1}\widetilde{n}_{j}^{\alpha}$ we obtain by induction
$\widetilde{n}_{j}\ \leq\ (2C_{1})^{1+\alpha+\cdots+\alpha^{j-1}}\,n^{\alpha^{j}}\ \leq\ (2C_{1})^{1/(1-\alpha)}\,n^{\alpha^{j}}.$
Therefore
$\prod_{j=0}^{t-1}\widetilde{n}_{j}^{\alpha}\leq(2C_{1})^{\alpha t/(1-\alpha)}\,n^{\alpha+\alpha^{2}+\cdots+\alpha^{t}}\leq(2C_{1})^{\alpha t/(1-\alpha)}\,n^{\alpha/(1-\alpha)}.$
Combining the previous estimates yields
$a_{n}\ \leq\ C_{0}\,C_{1}^{\,t}\,(2C_{1})^{\alpha t/(1-\alpha)}\,n^{\alpha/(1-\alpha)}.$
It remains to bound $t$. Write $x_{j}:=\log\widetilde{n}_{j}$. Since
$\widetilde{n}_{j+1}\leq 2C_{1}\widetilde{n}_{j}^{\alpha},$
we have
$x_{j+1}\leq\alpha x_{j}+\log(2C_{1}).$
Iterating this inequality gives
$x_{j}\leq\alpha^{j}\log n+\frac{\log(2C_{1})}{1-\alpha}.$
By the choice of $N_{0}$,
$\frac{\log(2C_{1})}{1-\alpha}\leq\frac{1}{2}\log N_{0}.$
Hence, whenever
$\alpha^{j}\log n\leq\frac{1}{2}\log N_{0},$
we obtain $x_{j}\leq\log N_{0}$, and therefore $\widetilde{n}_{j}\leq N_{0}$. Thus $t$ is at most the least integer $j$ such that
$\alpha^{j}\log n\leq\frac{1}{2}\log N_{0}.$
Equivalently, since $0<\alpha<1$,
$j\geq\frac{\log\log n-\log\bigl(\frac{1}{2}\log N_{0}\bigr)}{\log(1/\alpha)}.$
It follows that
$t\ll\log\log n.$
Consequently,
$C_{1}^{t}(2C_{1})^{\alpha t/(1-\alpha)}=(\log n)^{O(1)},$
and the claimed bound follows for all $n>N_{0}$. Enlarging $C_{2}$ covers the finitely many values $3\leq n\leq N_{0}$. ∎
We are now ready to present the following.
Proof of Theorem 1.6.
Fix $\varepsilon>0$. Choose $0<\eta<2506/4175$ so small that
$\frac{1}{(6681/4175-\eta)-1}\ \leq\ \frac{4175}{2506}+\frac{\varepsilon}{2}.$ (4.4)
Set $\delta=6681/4175-\eta$ and $\alpha=1/\delta\in(0,1)$. By Lemma 4.8, there exist constants $C_{1}\geq 2$ and $N_{0}$ such that
$a_{n}\leq C_{1}\,n^{\alpha}\,a_{\lceil C_{1}n^{\alpha}\rceil}\qquad(n\geq N_{0}).$
Applying Lemma 4.9 yields constants $B,C_{2}>0$ such that
$a_{n}\leq C_{2}\,n^{\alpha/(1-\alpha)}\,(\log n)^{B}\qquad(n\geq 3).$
Since $\alpha/(1-\alpha)=1/(\delta-1)$, the choice (4.4) gives
$n^{\alpha/(1-\alpha)}\ \leq\ n^{4175/2506+\varepsilon/2}.$
Finally, $(\log n)^{B}\leq n^{\varepsilon/2}$ for all sufficiently large $n$, so
$a_{n}\ \leq\ C_{\varepsilon}\,n^{4175/2506+\varepsilon}$
for all large $n$, and we adjust $C_{\varepsilon}$ to cover the finitely many smaller $n$. ∎
5. Concluding remarks
For the classical Hofstadter consecutive-sum sequence (1.1), the present paper provides both qualitative and quantitative progress on Problem 1.1. On the qualitative side, Corollary 1.5 shows that
$a_{n}=n+\omega(1),$
so the sequence omits infinitely many positive integers. On the quantitative side, Theorem 1.6 proves that for every $\varepsilon>0$,
$a_{n}\ll_{\varepsilon}n^{4175/2506+\varepsilon}.$
In particular, the sequence (1.1) has polynomial growth.
There remains, however, a substantial gap between these two conclusions, and the upper bound above is unlikely to be optimal. It is worth noting that, assuming the conjectural near-quadratic lower bound $|A-A|\gg|A|^{2-o(1)}$ for every finite convex set AA—a conjecture often attributed to Erdős and Hegyvári (see [3, p. 2])—the method of Section 4 would yield $a_{n}\leq n^{1+o(1)}$. However, even under this conjectural input, the present iteration argument still does not seem to give the stronger estimate $a_{n}\ll n$.
The available numerical evidence suggests that the true growth is much closer to linear. Writing
$b_{n}:=a_{n}-n,$
it is natural to ask for the true order of magnitude of $b_{n}$. To this end, we computed $b_{n}$ for $1\leq n\leq 30000$ and plotted the normalized quantities $b_{n}/n^{1/k}$ for $k=5,4,3,2$; see Figure 1. The ratio $b_{n}/n^{1/2}$ appears to decrease with $n$, whereas $b_{n}/n^{1/5}$ appears to increase, which is consistent with a power-law heuristic of the form
$b_{n}=\Theta\bigl(n^{\alpha+o(1)}\bigr)\qquad(n\to\infty)$
for some exponent $\alpha$ satisfying $1/5<\alpha<1/2$. Among these plots, the curve $b_{n}/n^{1/3}$ appears comparatively flat on the computed range, suggesting that $\alpha$ might be close to $1/3$, although it is not clear how representative this behavior is asymptotically.
Figure 1. Numerical data for sequence (1.1): the top panel shows $b_{n}=a_{n}-n$ (on a logarithmic $n$-axis), and the remaining panels plot the ratios $b_{n}/n^{1/k}$ for $k=5,4,3,2$ over $1\leq n\leq 30000$.
Acknowledgments. An earlier version of this paper, containing only the result $a_{n}=n+\omega(1)$, appeared as a short note in the comments section of the Erdős Problems website [2]. Matthew Bolan also independently obtained the qualitative result that the classical sequence omits infinitely many positive integers; see again the comments section of the Erdős Problems website [2]. We thank Ingo Althöfer for suggesting the generalization of Conjecture 1.2 to the setting of Definition 1.3. We also thank Boris Alexeev and Wolfram Bernhardt for computing and sharing the numerical data posted in the comments section of the Erdős Problems website [2].
Declaration of AI usage. An AI assistant (ChatGPT, model: GPT-5.2 Thinking) was used only during the brainstorming phase related to Theorem 1.6, and to draft the Python code used to generate Figure 1. Apart from this limited assistance, all mathematical arguments, proofs, and the writing of this paper were carried out by the author.

### 23 Mar 2026 v2

@@
 In this paper we prove that
-$n+\omega(1)\leq a_{n}\ll n^{4175/2506+o(1)}.$
+$n+\Omega(\log\log n)\leq a_{n}\ll n^{4175/2506+o(1)}.$
 In particular, $(a_{n})_{n\geq 1}$ omits infinitely many positive integers, thereby settling a conjecture from the OEIS entry A005243.
@@
 1. Introduction
-A self-generating greedy process, often discussed under the name Hofstadter sequences (see, for instance, [9] or [6, E31]), starts from a short initial list and then repeatedly adjoins sums of consecutive earlier terms.
-In its classical form, the seed is $(1,2)$, and one generates a strictly increasing sequence $(a_{n})_{n\geq 1}$ by setting $a_{1}=1$, $a_{2}=2$, and, for $k\geq 3$, taking $a_{k}$ to be the least integer greater than $a_{k-1}$ that can be written as a sum of at least two consecutive earlier terms:
+A self-generating greedy process, often discussed under the name Hofstadter sequence (see, for instance, [14] and [8, E31]), starts from a short initial list and repeatedly adjoins sums of consecutive earlier terms.
+In the classical case one begins with $(1,2)$ and defines a strictly increasing sequence $(a_{n})_{n\geq 1}$ by
+$a_{1}=1,\qquad a_{2}=2,$
+and, for $k\geq 3$, taking $a_{k}$ to be the least integer greater than $a_{k-1}$ that can be written as a sum of at least two consecutive earlier terms:
 $a_{k}=\sum_{i=p}^{q}a_{i}\qquad\text{for some }1\leq p\leq q\leq k-1,\ \ q-p\geq 1.$ (1.1)
-This is A005243 in the OEIS [8].
-The asymptotic behavior of $(a_{n})_{n\geq 1}$ was asked about by Hofstadter and has been recorded in several places, including [4, p.
-71], [5, p.
-83], [6, E31], and as Problem #423 on Bloom’s Erdős Problems website [2].
+This is sequence A005243 in the OEIS [11].
+The asymptotic behavior of $(a_{n})_{n\geq 1}$ was asked about by Hofstadter and has been recorded in several places, including [6, p.
+71], [7, p.
+83], [8, E31], and as Problem #423 on Bloom’s Erdős Problems website [5].
 Problem 1.1 ([6, 7, 8]).
@@
 The sequence (1.1) omits infinitely many positive integers.
-In this paper we prove Conjecture 1.2. In fact, our argument applies to a much broader class of related greedy sequences obtained from an arbitrary finite initial seed.
-Definition 1.3.
-Let $s\geq 2$ and let $\mathbf{u}=(u_{1},\dots,u_{s})$ be a seed of length $s$ consisting of positive integers (not necessarily distinct or ordered).
-Define a sequence of positive integers $(a_{n}^{\mathbf{u}})_{n\geq 1}$ by
-$a_{n}^{\mathbf{u}}:=u_{n}\qquad(1\leq n\leq s),$
-and for $n\geq s+1$ let $a_{n}^{\mathbf{u}}$ be the least integer greater than $a_{n-1}^{\mathbf{u}}$ that can be written as a sum of at least two consecutive earlier terms:
-$a_{n}^{\mathbf{u}}=\sum_{k=p}^{q}a_{k}^{\mathbf{u}}\qquad\text{for some }1\leq p\leq q\leq n-1,\ \ q-p\geq 1.$
-We also set
-$b_{n}^{\mathbf{u}}:=a_{n}^{\mathbf{u}}-n\qquad(n\geq 1).$
-This definition is well defined: for each $n\geq s+1$, the sum $a_{n-2}^{\mathbf{u}}+a_{n-1}^{\mathbf{u}}$ is a valid candidate for $a_{n}^{\mathbf{u}}$, and it is strictly greater than $a_{n-1}^{\mathbf{u}}$ because all terms are positive.
-Therefore the minimum in the definition of $a_{n}^{\mathbf{u}}$ always exists.
-In particular, we have $a_{n+1}^{\mathbf{u}}>a_{n}^{\mathbf{u}}$ for all $n\geq s$.
-When $\mathbf{u}=(1,2)$, the sequence $(a_{n}^{\mathbf{u}})_{n\geq 1}$ coincides with the classical sequence (1.1).
+For later use, write
+$b_{n}:=a_{n}-n\qquad(n\geq 1).$
+Our first result settles Conjecture 1.2.
 Theorem 1.3.
-For every integer $s\geq 2$ and every seed $\mathbf{u}$ of length $s$ consisting of positive integers, the sequence $(b_{n}^{\mathbf{u}})_{n\geq 1}$ is eventually nondecreasing; indeed, it is nondecreasing for all $n\geq s$.
-Moreover, $(b_{n}^{\mathbf{u}})_{n\geq 1}$ is unbounded.
-The quantity $b_{n}^{\mathbf{u}}=a_{n}^{\mathbf{u}}-n$ measures the excess of the current term over its index.
-For all sufficiently large $n$, it gives a lower bound for the number of omitted positive integers in $1,a_{n}^{\mathbf{u}}]$.
-As a consequence, we obtain the following corollary.
-Corollary 1.5.
-For every integer $s\geq 2$ and every seed $\mathbf{u}$ of length $s$ consisting of positive integers,
-$a_{n}^{\mathbf{u}}=n+\omega(1).$
-In particular, the set $\{a_{n}^{\mathbf{u}}:n\geq 1\}$ omits infinitely many positive integers.
-In particular, taking $\mathbf{u}=(1,2)$ in Corollary 1.5 proves Conjecture 1.2.
-For the classical sequence (1.1), we also obtain a quantitative upper bound. In fact, the same argument extends to every fixed positive seed; see Remark 4.1.
+The sequence $(b_{n})_{n\geq 1}$ is nondecreasing and unbounded. Consequently,
+$a_{n}=n+\omega(1),$
+and the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers.
+In fact, building on Theorem 1.3, we prove the following quantitative lower bound.
+Theorem 1.4.
+We have
+$a_{n}\geq n+\frac{\log\log n}{\log 20}-O(1).$
+We also obtain a quantitative upper bound.
 Theorem 1.5.
@@
 The proof of Theorem 1.5 combines the greedy structure of the sequence with a recent lower bound for difference sets of finite convex sets.
-Thus, besides proving that the sequence omits infinitely many positive integers, we also obtain a first polynomial upper bound toward Problem 1.1.
+Thus, besides settling Conjecture 1.2, we also obtain a first polynomial upper bound toward Problem 1.1.
+For related work on consecutive sums in integer sequences, see Beker [1], Hegyvári [9], and Tao [13].
 Notation
 We use Vinogradov’s asymptotic notation.
-For functions $f=f(n)$ and $g=g(n)$, we write $f=O(g)$, $f\ll g$, or $g\gg f$ to mean that there exists a constant $C>0$ such that $|f(n)|\leq Cg(n)$ for all sufficiently large $n$.
-If the implicit constant is allowed to depend on one or more parameters $z_{1},\dots,z_{r}$, we indicate this by writing $f\ll_{z_{1},\dots,z_{r}}g$, $g\gg_{z_{1},\dots,z_{r}}f$ or $f=O_{z_{1},\dots,z_{r}}(g)$.
-We write $f\asymp g$ or $f=\Theta(g)$ to mean that $f\ll g$ and $g\ll f$, and write $f=o(g)$ to mean $f(n)/g(n)\to 0$ as $n\to\infty$.
-We write $f=\omega(g)$ to mean that $g=o(f)$, equivalently, $|f(n)|/g(n)\to\infty$ when $n\to\infty$, assuming $g(n)>0$ for all sufficiently large $n$.
-In particular, $\omega(1)$ denotes any quantity tending to infinity as $n\to\infty$.
+For functions $f=f(n)$ and $g=g(n)$, we write $f=O(g)$, $g=\Omega(f)$, $f\ll g$, or $g\gg f$ to mean that there exists a constant $C>0$ such that $|f(n)|\leq Cg(n)$ for all sufficiently large $n$.
+If the implicit constant is allowed to depend on one or more parameters $z_{1},\dots,z_{r}$, we indicate this by writing $f=O_{z_{1},\dots,z_{r}}(g)$, $g=\Omega_{z_{1},\dots,z_{r}}(f)$, $f\ll_{z_{1},\dots,z_{r}}g$ or $g\gg_{z_{1},\dots,z_{r}}f$.
+We write $f\asymp g$ or $f=\Theta(g)$ to mean that $f\ll g$ and $g\ll f$, and write $f=o(g)$ to mean that $f(n)/g(n)\to 0$ as $n\to\infty$.
+We write $f=\omega(g)$ to mean that $g=o(f)$; in particular, $\omega(1)$ denotes any quantity tending to infinity as $n\to\infty$.
 Paper organization
-Section 2 collects the preliminary ingredients, including the characterization of integers that can be written as sums of at least two consecutive positive integers and a Diophantine finiteness consequence of a result of Schinzel and Tijdeman.
-In Section 3 we prove Theorem 1.4 and Corollary 1.5.
-Section 4 is devoted to the proof of the polynomial upper bound in Theorem 1.6.
-Finally, Section 5 contains concluding remarks and numerical observations.
+Section 2 collects the two preliminary ingredients used later: the characterization of sums of consecutive positive integers and a Diophantine finiteness consequence of the theorem of Schinzel and Tijdeman.
+Section 3 proves Theorem 1.3.
+Section 4 establishes the quantitative lower bound in Theorem 1.4.
+Section 5 proves the polynomial upper bound in Theorem 1.5.
+Finally, Section 6 contains concluding remarks on extensions, later improvements, and open questions.
 2. Preliminaries
@@
 is a sum of $2m\geq 2$ consecutive positive integers. ∎
-Next we record the Diophantine ingredient. We shall use the following consequence of a result of Schinzel and Tijdeman [7, Corollary 1].
+Next we record the Diophantine ingredient. We shall use the following consequence of a result of Schinzel and Tijdeman [10, Corollary 1].
 Lemma 2.2 ([10]).
@@
 $v^{2}+v+E=2^{k}$
-has only finitely many integer solutions $(v,k)\in\mathbb{Z}\times\mathbb{Z}$ with $k>2$.
+has only finitely many integer solutions $(v,k)\in\mathbb{Z}^{2}$ with $k>2$.
 Proof.
@@
 3. A qualitative lower bound and infinitely many omissions
-Throughout this section, we fix a seed $\mathbf{u}$ of length $s\geq 2$, and write
-$a_{n}:=a_{n}^{\mathbf{u}},\qquad b_{n}:=b_{n}^{\mathbf{u}}\qquad(n\geq 1).$
-Our goal is to show that $(b_{n})$ is unbounded. The first step is a simple monotonicity observation, which will later force eventual linearity if $(b_{n})$ were bounded.
+Throughout this section we write
+$b_{n}:=a_{n}-n\qquad(n\geq 1).$
+Our goal is to prove Theorem 1.3. The first step is a simple monotonicity observation.
 3.1. Monotonicity and eventual linearity
 Lemma 3.1.
-The sequence $(b_{n})_{n\geq s}$ is nondecreasing.
+The sequence $(b_{n})_{n\geq 1}$ is nondecreasing.
 Proof.
-For every $n\geq s$, the greedy construction gives $a_{n+1}>a_{n}$, hence $a_{n+1}\geq a_{n}+1$. Therefore, for $n\geq s$,
+For every $n\geq 1$, the greedy construction gives $a_{n+1}>a_{n}$, hence $a_{n+1}\geq a_{n}+1$. Therefore
 $b_{n+1}-b_{n}=(a_{n+1}-(n+1))-(a_{n}-n)=a_{n+1}-a_{n}-1\geq 0.\qed$
 Lemma 3.2.
-If $(b_{n})_{n\geq s}$ is bounded above, then there exist integers $n_{0}\geq s$ and $B$ such that
-$a_{n}=n+B\qquad\text{for all }n\geq n_{0}.$
+If $(b_{n})_{n\geq 1}$ is bounded above, then there exist integers $n_{0}\geq 2$ and $B$ such that
+$a_{n}=n+B\qquad(n\geq n_{0}).$
 Proof.
-By Lemma 3.1, the sequence $(b_{n})_{n\geq s}$ is a nondecreasing sequence of integers.
-If $(b_{n})_{n\geq s}$ is bounded above, then $(b_{n})_{n\geq s}$ must be eventually constant: there exist $n_{0}\geq s$ and an integer $B$ such that $b_{n}=B$ for all $n\geq n_{0}$.
+By Lemma 3.1, the sequence $(b_{n})_{n\geq 1}$ is a nondecreasing sequence of integers.
+If it is bounded above, then it is eventually constant: there exist $n_{0}\geq 2$ and an integer $B$ such that $b_{n}=B$ for all $n\geq n_{0}$.
 Hence $a_{n}=n+b_{n}=n+B$ for all $n\geq n_{0}$.
 ∎
-From now on we assume, for contradiction, that $(b_{n})_{n\geq s}$ is bounded above and hence that
+From now on we assume, for contradiction, that $(b_{n})_{n\geq 1}$ is bounded above and hence that
 $a_{n}=n+B\qquad(n\geq n_{0})$ (3.1)
-for some integers $n_{0}\geq s$ and $B$.
+for some integers $n_{0}\geq 2$ and $B$.
 3.2. Consecutive-block decompositions beyond the linear tail
@@
 $S_{n_{0}-1}:=\sum_{i=1}^{n_{0}-1}a_{i},\qquad\mathcal{C}:=\Bigl\{\sum_{k=p}^{n_{0}-1}a_{k}:\ 1\leq p\leq n_{0}\Bigr\}.$
-(Here $p=n_{0}$ contributes the empty sum $0\in\mathcal{C}$.) The next lemma isolates the only two ways a large term can be written as a sum of consecutive earlier terms once the tail is linear.
+(Here $p=n_{0}$ contributes the empty sum $0\in\mathcal{C}$.) The next lemma isolates the only two possible shapes of a large consecutive-block representation once the tail is linear.
 Lemma 3.3.
@@
 Assume (3.1). For every integer $r$ with
-$2^{r}\geq\max\{T,\ S_{n_{0}-1}+1,\ B+s+1\},$ (3.2)
+$2^{r}\geq\max\{T,\ S_{n_{0}-1}+1,\ B+3\},$ (3.2)
 there exist $C\in\mathcal{C}$ and an integer $v\geq T$ such that
@@
 Then $n\geq n_{0}$ and (3.1) gives $a_{n}=n+B=2^{r}$.
-Moreover, $n\geq s+1$, so by construction of the sequence there exist indices $1\leq p\leq q\leq n-1$ with $q-p\geq 1$ and
+Moreover, $n\geq 3$, so by construction of the sequence there exist indices $1\leq p\leq q\leq n-1$ with $q-p\geq 1$ and
 $2^{r}=a_{n}=\sum_{k=p}^{q}a_{k}.$
@@
 Proof of Theorem 1.3.
-Lemma 3.1 shows that $(b_{n})$ is nondecreasing for all $n\geq s$. It remains to prove that it is unbounded.
+Lemma 3.1 shows that $(b_{n})$ is nondecreasing. It remains to prove that it is unbounded.
 Assume, for contradiction, that $(b_{n})$ is bounded above.
 Then (3.1) holds by Lemma 3.2.
-Lemma 3.4 shows that for every sufficiently large $r$ we can find $C\in\mathcal{C}$ and $v\geq T$ such that (3.3) holds.
+By Lemma 3.4, for every sufficiently large $r$ there exist $C\in\mathcal{C}$ and $v\geq T$ such that (3.3) holds.
 The set
 $\mathcal{E}:=\{E_{C}:\ C\in\mathcal{C}\}$
-is finite. Hence by the pigeonhole principle there exists some $E\in\mathcal{E}$ for which the equation
+is finite. Hence, by the pigeonhole principle, there exists some $E\in\mathcal{E}$ for which the equation
 $v^{2}+v+E=2^{k}$
-has infinitely many integer solutions $(v,k)\in\mathbb{Z}\times\mathbb{Z}$ with $k>2$. This contradicts Corollary 2.3. Therefore $(b_{n})$ is unbounded, as claimed. ∎
-Proof of Corollary 1.5.
-By Theorem 1.4, the sequence $(b_{n})$ is nondecreasing for all $n\geq s$ and unbounded, hence $b_{n}\to\infty$. Therefore
+has infinitely many integer solutions $(v,k)\in\mathbb{Z}^{2}$ with $k>2$. This contradicts Corollary 2.3. Therefore $(b_{n})$ is unbounded.
+Since $(b_{n})$ is a nondecreasing unbounded sequence of integers, we have $b_{n}\to\infty$, and hence
 $a_{n}=n+b_{n}=n+\omega(1).$
-Let $M:=\max_{1\leq i\leq s}a_{i}$.
-For all sufficiently large $n$, we have $n\geq s$ and $a_{n}\geq M$.
-Hence every value attained by the sequence in $1,a_{n}]$ occurs among the first $n$ terms: indeed, the initial seed values are all $\leq M\leq a_{n}$, while $a_{m}>a_{n}$ for every $m>n$ because the sequence is strictly increasing from index $s$ onward.
-The set of distinct values among $a_{1},\dots,a_{n}$ therefore has cardinality at most $n$, so the number of positive integers in $1,a_{n}]$ omitted by the sequence is at least
-$a_{n}-n=b_{n}.$
-Since $b_{n}\to\infty$, the number of omitted integers up to $a_{n}$ is unbounded. In particular, the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers. ∎
-4. A quantitative upper bound for the classical sequence
-In this section we restrict attention to the classical sequence (1.1) and prove the bound
-$a_{n}\ll_{\varepsilon}n^{4175/2506+\varepsilon}.$
-Remark 4.1.
-For every fixed seed $\mathbf{u}$ of length $s\geq 2$ consisting of positive integers, the same exponent can also be obtained for $a_{n}^{\mathbf{u}}$, namely $a_{n}^{\mathbf{u}}\ll_{\mathbf{u},\varepsilon}n^{4175/2506+\varepsilon}$.
-Indeed, by Definition 1.3, the sequence $(a_{n}^{\mathbf{u}})_{n\geq s}$ is strictly increasing, so one may apply the argument of the present section to the tail
-$a_{s}^{\mathbf{u}},a_{s+1}^{\mathbf{u}},a_{s+2}^{\mathbf{u}},\dots$
-after making the obvious notational changes in the definitions of the relevant prefix sums and representable integers.
-Since this introduces additional bookkeeping but no essentially new idea, we present the details only for the classical seed $(1,2)$.
+Finally, because the sequence $(a_{n})$ is strictly increasing, exactly $n$ elements of $\{a_{m}:m\geq 1\}$ lie in the interval $1,a_{n}]$, namely $a_{1},\dots,a_{n}$. Therefore at least
+$a_{n}-n=b_{n}$
+positive integers in $1,a_{n}]$ are omitted by the sequence. As $b_{n}\to\infty$, the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers. ∎
+4. A quantitative lower bound for the classical sequence
+The following explicit consequence of Beukers’ work [2, Corollary 1] will be used.
+Lemma 4.1 ([2]).
+Let $D\in\mathbb{Z}\setminus\{0\}$. Every integer solution $(x,m)\in\mathbb{Z}^{2}$ of
+$x^{2}+D=2^{m}$
+satisfies
+$m<435+\frac{10\log|D|}{\log 2}.$
+Lemma 4.2.
+Let $\widehat{B}\geq 1$, and suppose that
+$b_{n}=\widehat{B}\qquad(n_{1}\leq n\leq n_{2})$
+for some integers $2\leq n_{1}\leq n_{2}$. Set
+$\widehat{T}:=n_{1}+\widehat{B},\qquad\widehat{S}:=\sum_{i=1}^{n_{1}-1}a_{i}.$
+If $r\in\mathbb{Z}$ satisfies
+$\max\{\widehat{T},\ \widehat{S}+1,\ \widehat{B}+3\}\leq 2^{r}\leq n_{2}+\widehat{B}$, (4.1)
+then there exist integers $x$ and $D\neq 0$ such that
+$2^{r+3}=x^{2}+D\qquad\text{and}\qquad|D|\leq 13\widehat{T}^{2}.$
+Proof.
+Since $(b_{n})_{n\geq 1}$ is nondecreasing by Lemma 3.1 and $b_{n_{1}}=\widehat{B}$, we have
+$b_{i}\leq\widehat{B}\qquad(1\leq i<n_{1}).$
+Hence
+$\widehat{S}=\sum_{i=1}^{n_{1}-1}a_{i}=\sum_{i=1}^{n_{1}-1}(i+b_{i})\leq\sum_{i=1}^{n_{1}-1}(i+\widehat{B})\leq\widehat{T}^{2}.$
+Now fix $r$ satisfying (4.1), and set $\widehat{n}:=2^{r}-\widehat{B}$. Then
+$n_{1}\leq\widehat{n}\leq n_{2},$
+so $a_{\widehat{n}}=\widehat{n}+\widehat{B}=2^{r}$. Also $\widehat{n}\geq 3$, and therefore the greedy definition of the classical sequence (1.1) yields indices $1\leq p<q\leq\widehat{n}-1$ such that
+$2^{r}=a_{\widehat{n}}=\sum_{k=p}^{q}a_{k}.$
+Since $2^{r}>\widehat{S}$, we must have $q\geq n_{1}$. If $p\geq n_{1}$, then
+$2^{r}=\sum_{k=p}^{q}(k+\widehat{B})=\sum_{j=p+\widehat{B}}^{q+\widehat{B}}j$
+is a sum of at least two consecutive positive integers, contradicting Lemma 2.1. Hence $p<n_{1}\leq q$.
+Set
+$\widehat{C}:=\sum_{k=p}^{n_{1}-1}a_{k},\qquad\widehat{v}:=q+\widehat{B}$.
+Then $0\leq\widehat{C}\leq\widehat{S}$ and
+$2^{r}=\widehat{C}+\sum_{j=\widehat{T}}^{\widehat{v}}j=\widehat{C}+\frac{\widehat{v}(\widehat{v}+1)-\widehat{T}(\widehat{T}-1)}{2}.$
+Thus
+$2^{r+1}=\widehat{v}^{2}+\widehat{v}+2\widehat{C}-\widehat{T}(\widehat{T}-1),$
+and therefore
+$2^{r+3}=(2\widehat{v}+1)^{2}+D,\qquad D:=8\widehat{C}-4\widehat{T}(\widehat{T}-1)-1.$
+Since $D\equiv-1\pmod{4}$, we have $D\neq 0$. Finally, using $0\leq\widehat{C}\leq\widehat{S}\leq\widehat{T}^{2}$, we obtain
+$|D|\leq 8\widehat{C}+4\widehat{T}(\widehat{T}-1)+1\leq 8\widehat{T}^{2}+4\widehat{T}^{2}+1\leq 13\widehat{T}^{2}.$
+This proves the lemma. ∎
+We are now ready to prove the quantitative lower bound for the sequence (1.1).
+Proof of Theorem 1.4.
+By Theorem 1.3, the sequence $(b_{n})$ is nondecreasing and unbounded. Hence, for each integer $\widehat{B}\geq 0$,
+$N(\widehat{B}):=\max\{n\geq 1:\ b_{n}\leq\widehat{B}\}$
+is well defined and finite. Set
+$M(\widehat{B}):=N(\widehat{B})+\widehat{B}+2\qquad(\widehat{B}\geq 0).$
+We claim that there exists an absolute constant $K\geq 2$ such that
+$M(\widehat{B})\leq K\,M(\widehat{B}-1)^{20}\qquad(\widehat{B}\geq 1).$ (4.2)
+Fix $\widehat{B}\geq 1$. If $N(\widehat{B})=N(\widehat{B}-1)$, then
+$M(\widehat{B})=N(\widehat{B})+\widehat{B}+2=N(\widehat{B}-1)+\widehat{B}+2=M(\widehat{B}-1)+1.$
+Since $M(\widehat{B}-1)\geq 2$, it follows that
+$M(\widehat{B})\leq M(\widehat{B}-1)^{20}.$
+Suppose now that $N(\widehat{B})>N(\widehat{B}-1)$. Since $(b_{n})$ is integer-valued and nondecreasing, we then have
+[…]
+Write
+$n_{1}:=N(\widehat{B}-1)+1,\qquad n_{2}:=N(\widehat{B}),\qquad\widehat{T}:=n_{1}+\widehat{B}$.
+Then
+$\widehat{T}=N(\widehat{B}-1)+1+\widehat{B}=N(\widehat{B}-1)+(\widehat{B}-1)+2=M(\widehat{B}-1).$
+Let
+$K_{0}:=2^{432}13^{10}.$
+We shall show that
+$n_{2}+\widehat{B}\leq 2K_{0}\widehat{T}^{20}.$
+Assume to the contrary that
+$n_{2}+\widehat{B}>2K_{0}\widehat{T}^{20}.$
+Then there exists an integer $r$ such that
+$K_{0}\widehat{T}^{20}<2^{r}\leq 2K_{0}\widehat{T}^{20}<n_{2}+\widehat{B}$.
+We now verify the hypotheses of Lemma 4.2. Since $\widehat{T}\geq 2$, we have
+$2^{r}>K_{0}\widehat{T}^{20}\geq\widehat{T}^{20}>\widehat{T}^{2}.$
+Moreover, for every $1\leq i\leq n_{1}-1$ we have $b_{i}\leq\widehat{B}-1$, because $n_{1}=N(\widehat{B}-1)+1$. Hence
+$a_{i}=i+b_{i}\leq i+\widehat{B}-1\qquad(1\leq i\leq n_{1}-1),$
+and therefore
+$\sum_{i=1}^{n_{1}-1}a_{i}\leq\sum_{i=1}^{n_{1}-1}(i+\widehat{B}-1)=\frac{(n_{1}-1)(n_{1}+2\widehat{B}-2)}{2}.$
+On the other hand,
+$\widehat{T}^{2}=(n_{1}+\widehat{B})^{2},$
+and a direct calculation shows that
+$2\widehat{T}^{2}-(n_{1}-1)(n_{1}+2\widehat{B}-2)=n_{1}^{2}+2n_{1}\widehat{B}+2\widehat{B}^{2}+3n_{1}+2\widehat{B}-2>0.$
+Thus
+$\sum_{i=1}^{n_{1}-1}a_{i}\leq\widehat{T}^{2}<2^{r}.$
+Also, from $2^{r}>\widehat{T}^{2}$ and $\widehat{T}\geq 2$, we obtain
+$2^{r}\geq\widehat{T}$.
+Finally, since $n_{1}\geq 2$, we have
+$\widehat{T}=n_{1}+\widehat{B}\geq\widehat{B}+2,$
+and therefore
+$2^{r}>\widehat{T}^{2}\geq(\widehat{B}+2)^{2}>\widehat{B}+3,$
+so in particular
+$2^{r}\geq\widehat{B}+3.$
+Hence Lemma 4.2 applies, and yields integers $x$ and $D\neq 0$ such that
+$2^{r+3}=x^{2}+D\qquad\text{and}\qquad|D|\leq 13\widehat{T}^{2}.$
+Applying Lemma 4.1, we obtain
+$r+3<435+\frac{10\log|D|}{\log 2}\leq 435+\frac{10\log(13\widehat{T}^{2})}{\log 2}.$
+Exponentiating, we get
+$2^{r}<2^{432}13^{10}\widehat{T}^{20}=K_{0}\widehat{T}^{20},$
+contrary to the choice of $r$. This contradiction proves that
+$n_{2}+\widehat{B}\leq 2K_{0}\widehat{T}^{20}.$
+[…]
+$N(\widehat{B})+\widehat{B}\leq 2K_{0}\,M(\widehat{B}-1)^{20},$
+and hence
+$M(\widehat{B})=N(\widehat{B})+\widehat{B}+2\leq(2K_{0}+2)\,M(\widehat{B}-1)^{20}.$
+Thus (4.2) holds with
+$K:=2K_{0}+2.$
+Taking logarithms in (4.2) and iterating, we obtain
+$\log M(\widehat{B})\leq 20^{\widehat{B}}\log M(0)+\bigl(1+20+\cdots+20^{\widehat{B}-1}\bigr)\log K$
+for all $\widehat{B}\geq 0$, whence
+$\log M(\widehat{B})\leq 20^{\widehat{B}}\log M(0)+\frac{20^{\widehat{B}}-1}{19}\log K\leq\widehat{A}\,20^{\widehat{B}}$,
+where
+$\widehat{A}:=\log M(0)+\frac{\log K}{19}.$
+[…]
+$\log n\leq\log M(\widehat{B})\leq\widehat{A}\,20^{\widehat{B}}$.
+Taking logarithms once more, we get
+$\widehat{B}=b_{n}\geq\frac{\log\log n-\log\widehat{A}}{\log 20}.$
+Finally, since $a_{n}=n+b_{n}$, it follows that
+$a_{n}\geq n+\frac{\log\log n}{\log 20}-\frac{\log\widehat{A}}{\log 20}.$
+This proves the theorem. ∎
+5. A polynomial upper bound for the classical sequence
+In this section we prove Theorem 1.5. The argument combines the greedy ordering of the representable integers with a lower bound for difference sets of finite convex sets.
 5.1. Greedy enumeration of the representable integers
@@
 This will yield a lower bound for $|\mathcal{R}_{m}|$, and hence, via the greedy enumeration established above, a recursive upper bound for $a_{n}$.
-The key additive-combinatorial input is the following difference-set bound for finite convex sets, due to Bloom [1, Theorem 2].
+The key additive-combinatorial input is the following difference-set bound for finite convex sets, due to Bloom [3, Theorem 2].
 Theorem 5.5 ([3]).
@@
 6. Concluding remarks
-For the classical Hofstadter consecutive-sum sequence (1.1), the present paper provides both qualitative and quantitative progress on Problem 1.1. On the qualitative side, Corollary 1.5 shows that
-$a_{n}=n+\omega(1),$
-so the sequence omits infinitely many positive integers. On the quantitative side, Theorem 1.6 proves that for every $\varepsilon>0$,
-$a_{n}\ll_{\varepsilon}n^{4175/2506+\varepsilon}.$
-In particular, the sequence (1.1) has polynomial growth.
-There remains, however, a substantial gap between these two conclusions, and the upper bound above is unlikely to be optimal.
-It is worth noting that, assuming the conjectural near-quadratic lower bound $|A-A|\gg|A|^{2-o(1)}$ for every finite convex set AA—a conjecture often attributed to Erdős and Hegyvári (see [3, p.
-2])—the method of Section 4 would yield $a_{n}\leq n^{1+o(1)}$.
-However, even under this conjectural input, the present iteration argument still does not seem to give the stronger estimate $a_{n}\ll n$.
-The available numerical evidence suggests that the true growth is much closer to linear. Writing
+6.1. Generalized seeds of size $2$ or larger
+Although the body of this paper has focused exclusively on the classical seed $(1,2)$, the same circle of ideas applies much more generally. If one starts from any fixed positive seed
+$\mathbf{u}=(u_{1},\dots,u_{s})\qquad(s\geq 2)$
+and then follows the same greedy rule, the monotonicity argument of Section 3, together with the eventual-linearity contradiction based on powers of $2$, still yields the analogue of Theorem 1.3.
+Likewise, the convex-difference-set argument of Section 5 extends, after routine changes in notation, to give a polynomial upper bound of the same shape as in Theorem 1.5.
+We have chosen not to include the full details here in order to keep the paper focused on Hofstadter’s original sequence (1.1).
+6.2. Later upper-bound improvements
+After this paper was completed, Sothanaphan pointed out [^1] See the comments section of the Erdős Problems website [5]. that Cushman’s recent bound
+$|A-A|\gg_{\varepsilon}|A|^{8/5+1/3440-\varepsilon}$
+for finite convex sets can be combined with the argument of Section 5 to yield the improved estimate
+$a_{n}\ll_{\varepsilon}n^{688/413+\varepsilon}.$
+See [12, 4]. Since this refinement does not alter the underlying method, we have chosen to keep the original exposition.
+6.3. Open questions
+A substantial gap remains between Theorems 1.4 and 1.5. It seems likely that the true growth of $a_{n}$ is much closer to linear than our present upper bound indicates. Writing
 $b_{n}:=a_{n}-n,$
-it is natural to ask for the true order of magnitude of $b_{n}$.
-To this end, we computed $b_{n}$ for $1\leq n\leq 30000$ and plotted the normalized quantities $b_{n}/n^{1/k}$ for $k=5,4,3,2$; see Figure 1.
+a natural next step is to determine the true order of magnitude of $b_{n}$. Most notably, one may ask whether $b_{n}$ is sublinear.
+Problem 6.1.
+Is it true that
+$a_{n}=n+o(n)?$
+Equivalently, does $b_{n}=o(n)$ hold?
+To complement this problem, we computed $b_{n}$ for $1\leq n\leq 30000$ and plotted the normalized quantities $b_{n}/n^{1/k}$ for $k=5,4,3,2$; see Figure 1.
 The ratio $b_{n}/n^{1/2}$ appears to decrease with $n$, whereas $b_{n}/n^{1/5}$ appears to increase, which is consistent with a power-law heuristic of the form
@@
 for some exponent $\alpha$ satisfying $1/5<\alpha<1/2$.
-Among these plots, the curve $b_{n}/n^{1/3}$ appears comparatively flat on the computed range, suggesting that $\alpha$ might be close to $1/3$, although it is not clear how representative this behavior is asymptotically.
+Among these plots, the curve $b_{n}/n^{1/3}$ appears comparatively flat on the computed range, suggesting that $\alpha$ might be close to $1/3$, although the available data are far from decisive.
 Figure 1.
-Numerical data for sequence (1.1): the top panel shows $b_{n}=a_{n}-n$ (on a logarithmic $n$-axis), and the remaining panels plot the ratios $b_{n}/n^{1/k}$ for $k=5,4,3,2$ over $1\leq n\leq 30000$.
+Numerical data for the sequence (1.1): the top panel shows $b_{n}=a_{n}-n$ (on a logarithmic $n$-axis), and the remaining panels plot the ratios $b_{n}/n^{1/k}$ for $k=5,4,3,2$ over $1\leq n\leq 30000$.
 Acknowledgments.
-An earlier version of this paper, containing only the result $a_{n}=n+\omega(1)$, appeared as a short note in the comments section of the Erdős Problems website [2].
+An earlier version of this paper, containing only the qualitative conclusion $a_{n}=n+\omega(1)$, appeared as a short note in the comments section of the Erdős Problems website [5].
 Matthew Bolan also independently obtained the qualitative result that the classical sequence omits infinitely many positive integers; see again the comments section of the Erdős Problems website [5].
-We thank Ingo Althöfer for suggesting the generalization of Conjecture 1.2 to the setting of Definition 1.3.
-We also thank Boris Alexeev and Wolfram Bernhardt for computing and sharing the numerical data posted in the comments section of the Erdős Problems website [2].
+We are grateful to Ingo Althöfer for encouraging us to think about extensions to more general initial seeds.
+We thank Yann Bugeaud for pointing out that the method in our proof of $a_{n}=n+\omega(1)$ can in fact be strengthened to yield the bound of Theorem 1.4.
+We also thank Norbert Hegyvári for sharing the papers [1, 9, 13] on consecutive sums.
+Finally, we thank Boris Alexeev and Wolfram Bernhardt for computing and sharing the numerical data posted in the comments section of the Erdős Problems website [5].
 Declaration of AI usage.
-An AI assistant (ChatGPT, model: GPT-5.2 Thinking) was used only during the brainstorming phase related to Theorem 1.6, and to draft the Python code used to generate Figure 1.
+An AI assistant was used only during the brainstorming phase related to Theorem 1.5, and to draft the Python code used to generate Figure 1.
 Apart from this limited assistance, all mathematical arguments, proofs, and the writing of this paper were carried out by the author.